OpenStudy (anonymous):

f(x)=x−9/(7x+2). Find an equation for the tangent line to the graph of f at x=11. Tangent line: y =

ganeshie8 (ganeshie8):

where are you stuck at ?

OpenStudy (anonymous):

in the begining

ganeshie8 (ganeshie8):

what have you tried

ganeshie8 (ganeshie8):

what do you need to write an equation of line ?

OpenStudy (anonymous):

i don't know how to do it !

ganeshie8 (ganeshie8):

equation of line with slope \(\color{red}{m}\) and y intercept \(\color{red}{b}\) :\[y = \color{red}{m}x + \color{red}{b}\]

ganeshie8 (ganeshie8):

thats called slope-intercept form of line, seen that before ?

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

good, calculus gives you slope \(\color{red}{m}\) of that tangent line

OpenStudy (anonymous):

f(x)=(x-9)/(7x+2)

ganeshie8 (ganeshie8):

\[f(x)=\dfrac{x-9}{7x+2}\] like this ?

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

Alright, here is the juicy formula \[\text{slope} = \dfrac{d}{dx} f(x) \] *read it like below : slope of tangent line at x=11 is same as the derivative of function at x = 11

ganeshie8 (ganeshie8):

for slope, you find the derivative of f(x) and evaluate it at x=11

ganeshie8 (ganeshie8):

can you find the derivative of f(x) ?